OpenStudy (lovelyharmonics):

Express the complex number in trigonometric form. -4

hartnn (hartnn):

do you know the formulas to convert complex numbers to trigonometric form ?

OpenStudy (lovelyharmonics):

no .-.

hartnn (hartnn):

ok., we would need x and y \(\large r = \sqrt{x^2+y^2} \\ \large \theta = \tan^{-1}\dfrac{y}{x}\) so compare -4 with x+iy whats x =...? y=... ?

OpenStudy (lovelyharmonics):

how are you getting 3 numbers out of -4!? o.o i get that i=- so y is probably 4 but where are you getting x from?

hartnn (hartnn):

we need only 2 numbers x and y see,